Consider a set up that corresponds between alphabetic characters as follows

Assuming the encryption algorithm is C=11m+7 mod 26, where m is the

message and C is the ciphertext.

I. Encrypt the word “CYBERCRIME

CR,10Marks

II. Derive the decryption algorithm


1
Expert's answer
2021-06-08T17:27:32-0400

The given encryption algorithm is C=11m+7 mod 26

A-0, B-1, C-2,D-3,E-4,F-5,G-6,H-7,I-8,J-9,K-10,L-11,M-12,N-13,O-14,P-15,Q-16,R-17,S-18,T-19,U-20,V-21,W-22,X-23,Y-24,Z-25

Now, given word is CYBERCRIME

Now, encrypting the given terms-

c=2

"(11\times 2+7)mod\ 26=29\ mod\ 26" =3

C=D

Y=24

"(11\times 24+7)\ mod \ 26 = 271 \ mod \ 26 = 11"

y=L


b= 1

"(11\times1+7)\ mod \ 26=18"

B= S


E=4

"(11\times 4+7)\ mod \ 26 = 51 \ mod \ 26 = 25"

E= Z


R=17

"(11\times 17 + 7)\ mod \ 26 = 194 \ mod \ 26 = 12"

R = M,


c=2

"(11\times 2+7)mod\ 26=29\ mod\ 26" =3

C=D


R=17

"(11\times 17 + 7)\ mod \ 26 = 194 \ mod \ 26 = 12"

R = M,


"I=8"

"(11\times 8 + 7)\ mod \ 26 = 95 \ mod \ 26 = 17"

I=R

M=12

"(11\times 12 + 7)\ mod \ 26 = 139 \ mod \ 26 = 9"

M=J

E=4

"(11\times 4+7)\ mod \ 26 = 51 \ mod \ 26 = 25"

E= Z

Hence, the encrypted word is DLSZMDMRJZ

ii) The decryption algorithm is given below -


"C=(11m+7)\ mod\ 26"

"\Rightarrow 11m=C-7 \ mod \ 26"

"\Rightarrow m= 11^{-1}(C-7 \ mod \ 26)"

Inverse of 11 in mod 26 is 19 Because "11\times 19 \ mod\ 26= 1\ mod \ 26"

Therefore Decryption algorithm will be "m=19(C-7) \ mod \ 26"

"m=19 C - 3\ mod \ 26"